Nikhef velo group meeting

Europe/Amsterdam
Nikhef

Nikhef

Kazu Carvalho Akiba, Kristof de Bruijn (Van Swinderen Institute), Wouter Hulsbergen
    • 3:30 PM 3:50 PM
      Round table 20m
      • Velo reinstallation
      • RF box production
      • Upgrade II efforts
      • Module tests at Nikhef
      • Velo reinstallation: planned for Jan 29-31 (3 days). Yutaro, Marco, Espen and Kazu will travel to CERN. Freek will be standby. Effort at CERN is coordinated by Keiran and Victor.
         
      • RF box production: there was an accident with the milling machine just before the break: The head of the machine ran into a corner of the box when doing the alignment. There was visible damage on the (premachined) outside. The box was taken from the machine, taken from the mold and re-measured. There seems to be no noticable deformation on the inside. The next step is to put the box again on the mold etc, and continue the maching of the outside.

       

       

      • module testing in H026:
        • MD took data cold with source
        • still need more data warm with source: MD would like to use Thursday and Friday
        • MD would like to have another source: but what kind of source?
        • can we do tests with testpulses? not sure if we can take data over slow control with testpulses.
        • equalization using fast links is still under development
        • master student Alfonso is looking at pixel-by-pixel variations of old data; didn't yet look at new data
        • it is time to start writing things down. Perhaps Alfonso can help with that